### v1

@@ 4.4.2 Multimodal Performance on Downstream Classification
Concatenation MLP 76.31 $\pm$ 0.45 77.18 $\pm$ 0.82 93.51 $\pm$ 1.06 97.45 $\pm$ 0.19
ResNet 76.38 $\pm$ 0.60 74.15 $\pm$ 1.76 92.20 $\pm$ 0.92 97.25 $\pm$ 0.53
FT-Transformer 75.51 $\pm$ 0.27 75.74 $\pm$ 0.81 85.40 $\pm$ 2.33 96.10 $\pm$ 0.39
FT-Mamba 75.80 $\pm$ 0.24 76.49 $\pm$ 0.60 89.08 $\pm$ 1.06 97.41 $\pm$ 0.36
Cross-attention FT-Transformer 75.71 $\pm$ 0.39 75.64 $\pm$ 0.34 93.53 $\pm$ 0.45 97.30 $\pm$ 0.77
FT-Mamba 73.36 $\pm$ 0.30 73.48 $\pm$ 0.98 96.96 $\pm$ 0.52 98.15 $\pm$ 0.31
AMF Module MLP 76.63 $\pm$ 0.22 76.86 $\pm$ 0.35 94.92 $\pm$ 0.43 97.88 $\pm$ 0.42
ResNet 76.50 $\pm$ 0.83 75.46 $\pm$ 0.92 95.09 $\pm$ 0.54 97.28 $\pm$ 0.70
FT-Transformer 75.62 $\pm$ 0.49 76.46 $\pm$ 1.57 93.83 $\pm$ 0.90 99.23 $\pm$ 0.35
FT-Mamba 76.97 $\pm$ 0.75 78.36 $\pm$ 0.12 94.76 $\pm$ 0.45 98.89 $\pm$ 0.11
Supervised ResNet-50 58.29 $\pm$ 4.16 93.00 $\pm$ 1.05
  • 1
    Frozen strategy, where the parameters of pretrained encoders did not participate in gradient updates.
  • 2
    Trainable strategy, where all weights in the model were trainable.
In the concatenation-based fusion, models using FT-Transformer and FT-Mamba underperform those using MLP and ResNet on both datasets. This contrasts with the superior performance of sequence-based encoders in unimodal training and contrastive pretraining. The performance drop can be attributed to the severe dimensional discrepancy between tabular features generated by sequence models ($D_{tab}=64$) and image features ($D_{img}=2048$). Sequence-based tabular encoders produce more compressed and lower-dimensional representations than MLP and ResNet, placing the tabular modality at a particular disadvantage during naive concatenation, where no interaction mechanism exists to compensate for disparities in feature scale and informativeness.
In contrast, the AMF module mitigates this imbalance by modulating unimodal features to comparable dimensions, restoring the advantages of sequence models in multimodal fusion. Compared to concatenation, the AMF module significantly improves performance, with an AUC increase of up to 1.87 percentage points on the OL3I dataset (with FT-Mamba) and an accuracy improvement of up to 8.43 percentage points on the DVM dataset (with FT-Transformer). When paired with MLP or ResNet tabular encoders, the AMF module also consistently outperforms concatenation across most settings.
Comparing the AMF module with the cross-attention module, the AMF module demonstrates a clear advantage under limited data conditions on the OL3I dataset. With FT-Mamba as the tabular encoder, AMF improves the AUC by 3.61 and 4.88 percentage points under frozen and trainable strategies, respectively. On the DVM dataset, where data is abundant, AMF performs comparably to the cross-attention module, suggesting that both methods are capable of effective representation learning when sufficient training samples are available.
Further comparison across both datasets reveals that AMF’s advantage over cross-attention is more prominent under the trainable strategy. This indicates that AMF is more responsive to encoder updates and better leverages the evolving feature representations during training. In contrast, the cross-attention module, with its complex and conservative full-attention structure, tends to be more stable in scenarios with frozen encoders and ample data. However, under the trainable strategy, its increased parameter complexity imposes optimization challenges, which hinders its effectiveness even on the DVM dataset.
4.4.3 Computational Efficiency of Fusion Modules
To compare the computational efficiency of the cross-attention and the proposed AMF module, we analyzed their number of parameters, floating-point operations (FLOPs), and inference latency in Table 3. All measurements were conducted with a batch size of 1. The inference latency was measured on an NVIDIA RTX 4090 GPU, averaged over 100 runs.
Table 3: Comparison of parameter count, FLOPs, and inference latency of the cross-attention and AMF modules.
Fusion Module #Param (M) OL3I DVM
FLOPs(GMac) Latency (ms) FLOPs(GMac) Latency (ms)
Cross-attention 9.49 3.67 1.54 3.17 1.21
AMF Module 8.53 0.0085 0.12 0.0085 0.12
@@ 4.4.3 Computational Efficiency of Fusion Modules
  • 1 GMac = $10^{9}$ multiply–accumulate operations.
  • The input image resolutions are 512 $\times$ 512 for the OL3I dataset and 128 $\times$ 128 for the DVM dataset.
Notably, the cross-attention module exhibits significantly higher parameter counts, FLOPs, and slower inference speed compared to the proposed AMF module, indicating greater model capacity and computational complexity. While such high-capacity architectures may enhance performance on large-scale datasets, they are also more susceptible to overfitting in low-data regimes, as they tend to memorize spurious correlations. In contrast, the lightweight AMF module offers a more streamlined architecture that generalizes better under data-scarce scenarios.
Importantly, the computational cost of cross-attention increases quadratically with the spatial resolution of input images, since its attention mechanism operates over the entire image patch sequence. This sensitivity to image size can greatly amplify both FLOPs and inference time under high-resolution settings. In contrast, the AMF module avoids token-level pairwise interactions and instead performs interaction based on the overall confidence scores of the image and tabular modalities, rendering its complexity effectively independent of image dimensions. This advantage ensures the scalability and practicality of AMF in real-world applications involving high-resolution inputs.
4.5 Robustness to Incomplete Tabular Data
In the IG analysis of the supervised tabular models, age was identified as the most important feature in the OL3I dataset by FT-Mamba, FT-Transformer, and ResNet. Therefore, we masked the true age value, and a new $R_{conf}$ of 1.03 was established based on unimodal performance, as shown in Table A.1. The AUC of multimodal models using incomplete tabular data, along with their absolute decrease compared to the complete data condition, are shown in Table 4.
Table 4: AUCs of multimodal models on incomplete OL3I data and their absolute decrease compared to the complete setting.
@@ 4.5 Robustness to Incomplete Tabular Data
AUC (%) Drop AUC (%) Drop
Concatenation - 70.02 $\pm$ 1.19 5.78 71.31 $\pm$ 2.20 5.18
Cross-attention - 72.58 $\pm$ 1.59 0.78 72.91 $\pm$ 0.50 0.57
AMF Module 0.95 74.88 $\pm$ 0.31 2.09 73.82 $\pm$ 0.84 4.54
AMF Module 1.03 75.07 $\pm$ 0.56 1.90 73.89 $\pm$ 0.78 4.47
In terms of AUC performance, the AMF module achieved the best results under incomplete tabular data conditions when $R_{conf}=1.03$, reaching 75.07% and 73.89% under frozen and trainable strategies, respectively—substantially outperforming other fusion methods. Regarding performance degradation, the cross-attention module exhibited the smallest AUC drop, suggesting a certain degree of robustness to incomplete tabular features. However, this small drop largely stems from its lower baseline performance when the tabular data is complete. Despite the smaller degradation, its absolute performance remains clearly inferior to that of the AMF module, indicating suboptimal fusion effectiveness. By contrast, the concatenation strategy suffered from over 5 percentage points of AUC drop in both missing conditions, reflecting a lack of effective compensation mechanisms in the presence of missing tabular inputs and thus poor robustness.
Further examining the response mechanism, the AMF module mitigates the negative impact of missing tabular features by explicit confidence control. Specifically, as the confidence in the tabular modality declines with increased feature missingness, the AMF module compensates by increasing $R_{conf}$ from 0.95 to 1.03, thereby encouraging the model to rely more heavily on the image modality. In contrast, the robustness of the cross-attention module primarily arises from implicit pattern fitting over large-scale data, lacking interpretability and generalizability in low-resource or high-missingness settings.
4.6 Ablation Studies on the AMF Module Design
4.6.1 Ablation Study on Modality Confidence Ratio
Table 5: AUCs of AMF-MedIT with different modality confidence ratios on the OL3I dataset.
$\mathbf{R_{conf}}$ $\mathbf{L_{img}}$ $\mathbf{L_{tab}}$ AUC (%)
[…]
None 2048 64 75.80 $\pm$ 0.24 76.49 $\pm$ 0.60
1.00 1024 1024 76.62 $\pm$ 0.90 78.33 $\pm$ 0.15
0.97 1009 1039 76.04 $\pm$ 1.09 78.08 $\pm$ 0.27
0.951 998 1059 76.97 $\pm$ 0.75 78.36 $\pm$ 0.12
0.10 342 1706 76.83 $\pm$ 0.37 78.16 $\pm$ 0.36
  • 1
    The value selected for FT-Mamba based on unimodal performance.
As shown in Table 5, the highest AUCs were achieved when $R_{conf}=0.95$, which aligns with the confidence ratio estimated from the unimodal performance of FT-Mamba and enables a balanced, task-consistent integration of both modalities. In contrast, settings based on mismatched modality pairs (e.g., $R_{conf}=0.97$ or $1.0$) resulted in performance degradation, suggesting that accurate allocation of unimodal feature dimensions is essential to maintain a balanced modality contribution and achieve effective fusion. In more extreme cases (e.g., $R_{conf}=0.1$ or None), the AUCs still declined compared to the optimal configuration, reinforcing that the adjustment of unimodal features is not arbitrary but grounded in meaningful modality-specific information.
These results collectively highlight that the effectiveness of the modulation-based fusion depends critically on appropriately scaling unimodal representations according to their task relevance. Therefore, it is necessary to introduce a modality confidence ratio based on actual modality reliability during fusion, which offers a simple yet interpretable way to achieve balanced unimodal feature scaling.
4.6.2 Ablation Study on Loss Functions
$\mathcal{L}_{\text{ce}}$ $\mathcal{L}_{\text{leakage}}$ $\mathcal{L}_{\text{density}}$ AUC (%)
[…]
$\times$ $\times$ 73.11$\pm$3.99 78.07$\pm$0.64
$\times$ 75.46$\pm$1.71 78.05$\pm$0.36
$\times$ 74.11$\pm$3.49 78.08$\pm$0.35
76.97$\pm$0.75 78.36$\pm$0.12
@@ 4.6.2 Ablation Study on Loss Functions
The AUCs under different loss configurations are presented in Table 6. The best performance is achieved when all three losses are jointly applied. Removing the leakage loss $\mathcal{L}_{\text{leakage}}$ leads to a substantial drop in performance, especially under the frozen strategy (-2.86% AUC), indicating that the absence of information flow regulation across modalities causes significant semantic leakage and degraded fusion quality. Similarly, omitting the density loss also reduces performance and introduces instability, as it fails to constrain the feature magnitude alignment between the image and tabular modalities, which can impede effective fusion.
@@
The improvement brought by these two auxiliary losses is more evident in the frozen training setting, where the fusion module receives no gradient from the backbone. In such cases, relying solely on the classification loss $\mathcal{L}_{\text{ce}}$ is insufficient for guiding the training of the AMF module, leading to higher performance variance and lower overall accuracy. These results collectively validate the necessity of both density and leakage losses for effective multimodal training, as they enhance representation compatibility, stabilize optimization, and ultimately improve classification performance.
@@ 5 Discussion
In this work, we propose AMF-MedIT, an align-modulation-fusion framework for medical image and tabular data. Our motivation stems from the significant cross-modal differences in dimension and confidence levels during fusion, which are further exacerbated by the scarcity of medical data. To address this, we propose the AMF module, a simple yet efficient solution for multimodal fusion. The AMF module not only achieves effective multimodal classification performance on both medical and natural datasets, but also offers high computational efficiency and strong adaptability to incomplete tabular data. Furthermore, for the noisy medical tabular data, we present FT-Mamba for tabular feature extraction. It achieves a favorable balance between data efficiency and feature discriminability and plays a pivotal supervisory role in image-tabular contrastive pretraining.
5.1 Discussion of Fusion Methods
In the study of fusion methods, we compared our proposed AMF module with two widely adopted approaches—naive feature concatenation and the cross-attention mechanism—across three key aspects: performance, computational efficiency, and robustness to missing tabular data. Naive concatenation is a straightforward and commonly used approach in medical multimodal learning. However, its fixed and non-interactive network structure lacks the capacity to regulate the dimensionality and contribution of each modality. As a result, compact tabular features are placed at a disadvantage, leading to suboptimal fusion performance. Furthermore, the rigidity of this approach makes it less robust to variations in downstream data distributions, limiting its adaptability in real-world clinical scenarios.
The cross-attention module, on the other hand, takes multimodal token sequences with equal dimensions as input. By leveraging explicit token-level attention, it enables deeper cross-modal interaction and demonstrates strong robustness in handling incomplete tabular data. Nevertheless, its full attention architecture introduces high computational complexity and memory overhead, which may hinder its applicability in data-limited medical contexts or real-time clinical deployments.
To strike a better balance between effectiveness and efficiency, the AMF module replaces token-level attention with a lightweight, confidence-driven modulation strategy. Rather than enforcing equal or fixed feature dimensions, AMF adaptively adjusts the unimodal feature dimensions according to its global confidence. In addition, we propose a density loss that serves as a soft normalization mechanism to align the density of features across modalities. The adjustments in both dimension and density help mitigate scale mismatches between heterogeneous modalities and facilitate the formation of a unified cross-modal representation space. Moreover, our design incorporates a feature mask along with a leakage loss, which permits flexible information flow and compression, enabling the unified space to preserve complementary modality-specific information. This unified yet relaxed-constrained representation space reconciles the representational and confidence discrepancies between tabular and image modalities, supporting effective multimodal integration as confirmed by our experimental results. At the same time, the module’s simplicity allows for efficient fine-tuning with limited data and fast inference, making it practical for real-world deployment.
As another simplification, we introduce a scalar modality confidence ratio $R_{conf}$ which is derived from preliminary empirical results. The incorporation of prior knowledge is particularly valuable in the medical domain. On the one hand, it helps mitigate data scarcity by leveraging external expertise and reduces the reliance on purely data-driven learning. On the other hand, it enables the integration of medical knowledge or clinical experience into future automated systems, enhancing the interpretability of the fusion process and its acceptance among medical professionals. Beyond facilitating effective fusion, $R_{conf}$ also enhances the model’s adaptivity to various clinical scenarios, including tabular data missingness in our experiments, image quality degradation, and even modality absence. Importantly, AMF achieves this with fewer parameters and orders-of-magnitude lower computational cost than cross-attention, making it more suitable for data-scarce, latency-sensitive, or resource-constrained medical applications.
5.2 Discussion of Tabular Encoders for Medical Data
In the study of tabular encoders, we compared the performance of four representative architectures—MLP, ResNet, FT-Transformer, and FT-Mamba—and further conducted interpretability analyses to understand their working mechanisms and roles in contrastive learning. As a kind of structural text, medical tabular data also exhibits sequential characteristics. For such data, sequence models like FT-Mamba and FT-Transformer demonstrate superior feature discrimination through their respective selection and self-attention mechanisms. This ability to automatically identify relevant features from a large pool of redundant variables is particularly important in processing EHRs, which encompass a wide range of diseases, time spans, and heterogeneous variables. Enhanced feature discrimination not only reduces the need for manual feature engineering but also enables the model to focus on clinically relevant patterns, potentially uncovering latent medical associations. However, among these two sequence models, FT-Transformer’s reliance on full attention matrices leads to overfitting on small datasets, while FT-Mamba is more efficient in low-data settings by selectively focusing on relevant information.
Due to its advantages, FT-Mamba provides effective supervision for imaging modality during contrastive pretraining, guiding the encoder to flexibly focus on multiple key regions. In contrast, ResNet exhibits strong locality in tabular feature extraction, showing a marked focus on a single feature—age—which induces a singular and coarse focus in the image modality. The visual attention is primarily concentrated on the lumbar spine, likely because the correlation between patient age and bone density (Jiang et al., 2022) is implicitly leveraged by the model during contrastive learning. Our interpretability analysis reveals that tabular encoders with different representational capacities exert distinctly different supervisory effects on the image modality during contrastive learning. Therefore, the selection of the tabular encoder plays a crucial role in multimodal pretraining, as it directly influences the representational power of the learned features.
5.3 Limitations and Future Directions
Our study has several limitations. First, the results on the DVM dataset indicate that FT-Mamba still underperforms FT-Transformer on large-scale data. While FT-Mamba serves as a compromise for limited medical data, we expect FT-Transformer to perform better as large-scale medical datasets become more common. Second, the $R_{conf}$ configuration in the AMF module is sensitive, with slight variations significantly affecting performance in our experiments. Third, our experiments were limited to two multimodal datasets, and further validation on a broader range of medical datasets is needed. Finally, our model is restricted to classification tasks, and other medical image-tabular applications, such as segmentation and report generation, require further exploration.
Future research could investigate automated optimization of $R_{conf}$ based on dataset characteristics and develop guidelines informed by domain experts to improve its robustness. Additionally, evaluating our approach on diverse datasets, including those from different imaging modalities like magnetic resonance imaging and ultrasound, would enhance its generalizability. Extending the AMF module to support segmentation tasks could also broaden its applicability in medical research.

### v2

@@ 4.5.2 Ablation Study on Multimodal Fusion Module
Concact [6] - 75.80 $\pm$ 0.24 76.49 $\pm$ 0.60 - - 89.08 $\pm$ 1.06 97.41 $\pm$ 0.36 - -
Cross-attn. [25] 9.49 73.36 $\pm$ 0.30 73.48 $\pm$ 0.98 3.67 1.54 96.96 $\pm$ 0.52 98.15 $\pm$ 0.31 3.17 1.21
AMF (ours) 8.53 76.97 $\pm$ 0.75 78.36 $\pm$ 0.12 0.0085 0.12 94.76 $\pm$ 0.45 98.89 $\pm$ 0.11 0.0085 0.12
@@
  • Note: FLOPs = floating-point operations (in GMac, $10^{9}$ multiply–accumulate operations). Lat. = latency (in ms).
On the OL3I dataset, AMF achieves the highest AUC (78.36%) under the trainable strategy, outperforming concatenation (76.49%) and cross-attention (73.48%). On the DVM dataset, AMF remains competitive under a trainable strategy (98.89%), demonstrating robust performance across datasets of different scales. These results confirm that simple fusion methods lack sufficient representational capacity, while highly complex methods like cross-attention are less robust under low-data regimes.
From an efficiency perspective, the cross-attention module incurs substantially higher parameters, FLOPs, and inference latency compared to AMF, reflecting its greater computational complexity. Notably, its cost grows quadratically with the spatial resolution of input images, due to token-level pairwise attention. In contrast, AMF performs interaction based on modality-level confidence scores rather than token-level pairwise computation, making its complexity effectively independent of image size. This design ensures that AMF is both scalable and practical, providing a strong balance between performance and efficiency for real-world multimodal applications.
4.5.3 Ablation Study on Loss Functions
To evaluate the contribution of the proposed auxiliary loss functions—magnitude loss and leakage loss, we fixed $r_{\text{conf}}=0.95$ and compared four different loss configurations during fine-tuning: (1) $\mathcal{L}_{\text{fine-tuning}}=\mathcal{L}_{\text{ce}}$; (2) $\mathcal{L}_{\text{fine-tuning}}=\mathcal{L}_{\text{ce}}+\lambda_{1}\mathcal{L}_{\text{leakage}}$; (3) $\mathcal{L}_{\text{fine-tuning}}=\mathcal{L}_{\text{ce}}+\lambda_{2}\mathcal{L}_{\text{magnitude}}$; (4) $\mathcal{L}_{\text{fine-tuning}}=\mathcal{L}_{\text{ce}}+\lambda_{1}\mathcal{L}_{\text{leakage}}+\lambda_{2}\mathcal{L}_{\text{magnitude}}$ (ours).
$\mathcal{L}_{\text{ce}}$ $\mathcal{L}_{\text{leakage}}$ $\mathcal{L}_{\text{magnitude}}$ OL3I AUC (%) DVM AUC (%)
@@ 4.5.3 Ablation Study on Loss Functions
$\times$ $\times$ 73.11 $\pm$ 3.99 78.07 $\pm$ 0.64 88.31 $\pm$ 1.07 97.30 $\pm$ 0.36
$\times$ 75.46 $\pm$ 1.71 78.05 $\pm$ 0.36 91.01 $\pm$ 0.55 96.79 $\pm$ 0.70
$\times$ 74.11 $\pm$ 3.49 78.08 $\pm$ 0.35 91.20 $\pm$ 0.57 97.08 $\pm$ 0.30
76.97 $\pm$ 0.75 78.36 $\pm$ 0.12 94.76 $\pm$ 0.45 98.89 $\pm$ 0.11
@@
The AUCs under different loss configurations are presented in Table 4. The best performance is achieved when all three losses are jointly applied. Removing the leakage loss $\mathcal{L}_{\text{leakage}}$ leads to a substantial drop in performance, especially under the frozen strategy (-2.86% AUC), indicating that the absence of information flow regulation across modalities causes significant semantic leakage and degraded fusion quality. Similarly, omitting the magnitude loss also reduces performance and introduces instability, as it fails to constrain the feature magnitude alignment between the image and tabular modalities, which can impede effective fusion.
@@
The improvement brought by these two auxiliary losses is more evident in the frozen training setting, where the fusion module receives no gradient from the backbone. In such cases, relying solely on the classification loss $\mathcal{L}_{\text{ce}}$ is insufficient for guiding the training of the AMF module, leading to higher performance variance and lower overall accuracy. These results collectively validate the necessity of both magnitude and leakage losses for effective multimodal training, as they enhance representation compatibility, stabilize optimization, and ultimately improve classification performance.
@@ 5 Discussion
In this work, we proposed AMF-MedIT, an efficient framework for integrating medical image and tabular data. Our approach was motivated by the substantial heterogeneity between modalities, particularly in terms of dimensionality and confidence levels, as well as the challenges posed by small and noisy clinical datasets. To address these issues, we designed a twofold strategy: (1) introducing prior knowledge through the AMF fusion module to achieve robust and adaptive integration across modalities, and (2) leveraging self-supervised contrastive pretraining together with an efficient tabular encoder, FT-Mamba, to enhance multimodal representation learning under limited data.
Compared with existing methods, AMF-MedIT demonstrates consistent advantages in both small-scale medical datasets and large-scale natural datasets. In clinical practice, multimodal learning often faces the dual challenges of data scarcity and noise. Medical image–tabular multimodal datasets remain limited in availability due to the high costs of data acquisition and the complexities of data management; moreover, missing clinical records or measurement errors can lead to unreliable tabular features, while artifacts and acquisition noise may compromise image quality. These challenges highlight that practical medical artificial intelligence (AI) systems require not only powerful models but also strategies that enhance efficiency and adaptability to limited data. Our experiments show that AMF-MedIT strikes an effective balance between performance and data efficiency, benefiting from its simple yet adaptive design. Specifically, instead of relying on cross-attention-based modules, we adopt a modulation-based fusion strategy that incorporates prior knowledge, thereby alleviating the training burden and avoiding the data inefficiency often caused by overly complex fusion modules. Furthermore, thanks to its modality-agnostic confidence adjustment mechanism, AMF-MedIT exhibits robust tolerance to noise in both image and tabular modalities. In addition, the introduction of feature masks and leakage loss enables adaptive adjustment to changing downstream data conditions without requiring structural modifications to the network, thereby providing broader clinical applicability compared with methods tailored to specific noise types.
Ablation studies further highlight the key contributions of both the FT-Mamba encoder and the AMF module. First, the tabular encoder plays a central role in maintaining both high accuracy and data efficiency. Interpretability analyses reveal that FT-Mamba captures fine-grained, discriminative tabular features more effectively than MLP and ResNet encoders, which are constrained by local attention. Moreover, by employing a selective mechanism, it achieves efficient feature extraction that remains robust under limited data. Compared with transformer-based models that rely on full attention, it is capable of filtering out irrelevant features while preserving medically meaningful signals, without overfitting to spurious noise patterns introduced by model complexity. This capability also enhances contrastive pretraining and guides the image encoder toward learning precise and clinically relevant attention patterns. Second, the AMF module achieves strong performance with remarkably low computational complexity, in contrast to concatenation- or attention-based approaches. Importantly, its complexity is independent of image resolution, which is advantageous for small datasets and satisfies the real-time requirements of clinical applications. Finally, loss function ablations confirm that leakage loss and density loss contribute to training stability and overall performance.
Despite the observed benefits, our study has several limitations. First, while FT-Mamba provides a good trade-off under limited data, it still underperforms FT-Transformer on large-scale datasets, suggesting that FT-Transformer may be preferable as larger medical datasets become available. Second, the configuration of the $r_{\text{conf}}$ parameter in the AMF module remains sensitive. Our experiments on missing tabular features indicate that estimates derived from frozen strategies may not generalize well to trainable settings. More precise tuning strategies, potentially incorporating expert knowledge, are needed to improve robustness. Third, our validation was restricted to two multimodal datasets, and further evaluation on a wider range of clinical data is necessary. Lastly, our framework was applied only to classification tasks; its potential for other medical applications, such as segmentation or automated report generation, remains unexplored.
Looking forward, several directions can be pursued. Future work could focus on developing automated procedures to optimize $r_{\text{conf}}$ based on dataset-specific characteristics, while integrating domain expertise to guide its estimation. Extending the evaluation to more diverse modalities, such as MRI and ultrasound, will further establish the generalizability of our framework. Moreover, expanding AMF-MedIT beyond classification to support segmentation and report generation tasks could broaden its impact on multimodal medical AI.